OpenStudy (anonymous):

Estimate the indicated derivative by any method. (Round your answer to three decimal places.) s = t − t2; estimate ds/dt t = 4

OpenStudy (anonymous):

*s=t-t^2

OpenStudy (anonymous):

what does "estimate" mean? the derivative of \(t-t^2\) is \(1-2t\)

OpenStudy (anonymous):

if you plug in 4 you get -7

OpenStudy (anonymous):

How did you know t-t^2 is 1-2t?

OpenStudy (anonymous):

i guess you didn't get to the part about the "power rule" yet the derivative of \(t\) is 1, because \(t\) is a line with slope 1 the derivative of \(t^2\) is \(2t\) by the all mighty power rule, but you can get it by hand if you like it will be a bit of work, but it is easy if you memorize that \[\frac{d}{dx}[x^n]=nx^{n-1}\]

OpenStudy (anonymous):

That's extremely helpful. Thanks so much!